OpenStudy (anonymous):

Using substitution, what is the solution of the system? 4y-2x=10 3y + x = 15

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you that was the part i was stuck on! but you still have to substitute 4 for y in one of the first problems I did 3y+x=15 3*4+x=15 12-12+x=15-12 x=3